OpenStudy (anonymous):

\[\int\limits_{}^{}\tan^3x secx dx \]

OpenStudy (raden):

first, change tan^3 x = tan^2x tanx

OpenStudy (anonymous):

I tried this \[\int\limits_{}^{}(\tan^2x. tanx. secx) dx \] \[\int\limits_{}^{}(\sec^2x-1) (tanx. secx) dx \]

OpenStudy (slaaibak):

Yeah, continue with that and set u = sec x du = sec x tan x

OpenStudy (slaaibak):

du = sec x tan x dx

OpenStudy (raden):

ok, then multiply of them by use distributive property = int (sec^2x secxtanx - secxtanx) dx = int (sec^2x secxtanx) dx - int (secxtanx) dx

OpenStudy (raden):

case I : int (sec^2x secxtanx) dx use int by subs, like slaaibak said : u=secx case II : int (secxtanx) = secx (i sure u have remembered) :p

OpenStudy (anonymous):

I think I do :) thanks @RadEn and @slaaibak